OpenStudy (anonymous):

Which ratio is proportional to 4/15 a.12/45 b.2/7 c.16/30 d.36/75

OpenStudy (anonymous):

A.12/45

OpenStudy (anonymous):

For something to be proportional, it means they are equal in some way, if you notice something about 4/15, well, you can multiply it by 3 over 3\[\frac{ 4 }{ 15 }=\frac{ 3*4 }{ 3*15 }=\frac{ 12 }{ 45 }\]you can do this because 3/3 is 1, so in the end you're not changing the actual fraction